1.Drive the base units of the following
a . pressure
b . pressure
1
Expert's answer
2021-07-13T10:42:23-0400
"[P]=\frac{[F]}{[A]}\\ [P]=\frac{[ma]}{[A]}\\ [P]=\frac{kgms^{-2}}{m^2}\\ [P]=\frac{kg}{ms^2}\\"